OpenStudy (anonymous):

The incidence of cystic fibrosis, a recessive genetic disorder in the Caucasian population of United States, is 1 in every 2,500 individuals. Find the number of heterozygous carriers. A.0.0392 B.0.0254 C.0.045

OpenStudy (aaronq):

The only equations you need for these types of problems: p + q =1 p = frequency of the dominant allele in the population q = frequency of the recessive allele in the population and p^2 + 2pq + q^2 =1 p^2 = percentage of homozygous dominant individuals q^2 = percentage of homozygous recessive individuals 2pq = percentage of heterozygous individuals in the question they tell you that 1/2500 have the disorder.. meaning q^2=1/2500 they're asking you to find 2pq go!